# Icy or Rocky? Convective or Stable?

### v1

@@ 3 Uranus
Figure 3: Density, pressure, and temperature profiles for the four Uranus models. Top: Density vs. normalised radius. For comparison, also shown are empirical solutions from Neuenschwander & Helled (2022) (grey area). The panel in the top right shows Uranus’ measured (French et al., 2024, black) and calculated (coloured) gravitational moments. We also show Uranus’ gravitational moments according to Jacobson & Park (2025) (grey) who found different results (COR vs. COO) corresponding to different ring centre definitions. The coloured uncertainties depict estimated errors from the Theory of Figures. Bottom: Temperature vs. pressure. Also shown is the predicted region of ionic water inferred by Redmer et al. (2011). The dashed lines show the demixing line, below which hydrogen, helium, and water are expected to become immiscible (Howard et al., 2025).
We present four Uranus models: U1, U2, U3, and U4. The inferred internal structures and compositions of the Uranus models are summarised in Figure 2. Models U1 and U3 correspond to water-rich interior solutions. Their respective rock-to-water ratios are low, just 0.04 and 0.08, consistent with Uranus being an ”ice giant”. These models also possess similar total H-He abundances (14% and 11% of Uranus’ mass, respectively). However, their convection and temperature profiles vary significantly. Model U1 is mostly convective, has four different convection zones and reaches a central temperature below 13,000 K. On the other hand, model U3 has two large stable regions. The innermost stable region covers more than half (54%) of Uranus’s radius and the central temperature is above 20,000 K.
@@
The density profiles and temperature-pressure profiles are presented in Figure 3. We find that the density profiles for models U2 and U4 are very similar. They both have a density discontinuity ($\Delta\rho>$ 1 g cm-3) at $\sim$ 75% of Uranus’ radius and reach central densities of $\sim$4.5 g cm-3. This density discontinuity around $3\cdot 10^{5}$ bar leads to a temperature jump that is 500 K higher for U2 compared to U4. Models U1 and U3 reach higher central densities (6.6 and 7.8 g cm-3, respectively) than U2 and U4. There are also no clear jumps in density in models U1 and U3. We further observe that all models leave the solution space discovered by Neuenschwander & Helled (2022) around a normalized radius of 0.8. This finding demonstrates how assumptions made by the modeller (such as the use of three consecutive polytropes in Neuenschwander & Helled (2022)) can result in biased outcomes and exclude parts of the solution space. In contrast, our randomized approach represents a notable improvement, expanding the scope of the solution space.
For comparison, we show the measured gravitational moments from Jacobson & Park (2025) (in grey) alongside those from French et al. (2024) (in black) in Figure 3. Jacobson & Park (2025) only reported formal (not realistic) uncertainties for their two results based on precession rates of either the ring centers of opacity (COO) or the geometric ring midlines (COR). They concluded that their knowledge of the gravitational moments is not better compared to that postulated by French et al. (2024).
@@ 4 Neptune
We present four Neptune models N1, N2, N3 and N4. Figure 4 shows sketches of the inferred structure and bulk composition of our Neptune models. Models N1 and N4 are consistent with the notion of Neptune being an ”ice giant”. They have water-dominated interiors with rock-to-water ratios of 0.20 and 0.26, respectively, and H-He mass fractions of 12% and 10%, respectively. However, there are also significant differences. Model N1 has a central density higher than 18 g cm-3, while the central density of model N4 is only 6.2 g cm-3. Although both models include stable regions, they are more prevalent for model N1. For example, model N4 has a fully convective deep interior while the innermost region of model N1 is stable against convection and has a composition gradient. Model N4 initially has a larger temperature gradient compared to model N1, leading to a first appearance of ionic water at relatively larger normalised radii (0.89 vs. 0.83) with higher temperatures (3100 K vs. 2500 K).
@@
Models N2 and N3 represent Neptune models that are rock-dominated, supporting the idea of Neptune being a ”rock giant” (for example Helled & Fortney, 2020; Teanby et al., 2020). Their rock-to-water ratios are found to be 1.78 and 1.72, with more than 50% of the planetary mass consisting of rocks. Although both models have large convective zones, the interior temperature profiles are rather different. Model N2 has a central temperature of 28,000 K, while the central temperature of model N3 is 12,000 K. Both of these inferred temperatures are significantly higher than the temperature inferred by standard adiabatic models which are of the order of 6000 K (for example Nettelmann et al., 2013). In addition, in model N3 H-He is found also in the planetary centre with 10 % of the composition by mass. We note that the existence of H-He in the deep interior is consistent with formation models of Neptune with composition gradients (for example Valletta & Helled, 2022). The central region in model N2, on the other hand, is enriched with iron (18 % by mass).
@@
The inferred density profiles and pressure-temperature profiles are shown in Figure 5. For Neptune, we find four very unique density profiles. Model N1 has the highest central density of 18.4 g cm-3. The density in model N2 is below 3 g cm-3 in the outer half of the planet until it reaches a large density discontinuity ($\Delta\rho>3$ g cm-3). Models N3 and N4 have similar central densities (5.8 and 6.2 g cm-3, respectively), but differ with respect to the location of their density discontinuity. The discontinuities are located at a normalised radius of $\sim$0.7 for model N3 and $\sim$0.6 for model N2. We further find that models N1, N2, and N4 have a temperature jump larger than 500 K between pressures of 200 and 2000 bars. Also in the case of Neptune, there are significant differences in the density profiles in comparison to the solution space found by Neuenschwander & Helled (2022). This is observed at multiple locations (normalised radii of around 0.05, 0.4, and 0.75) with the differences being more notable compared to Uranus. Our randomized approach indeed covers a wider space of solutions compared to parameterized approaches.
@@ 5.1 Bulk composition
The inferred bulk compositions of our Uranus and Neptune models are summarised in Figure 6. We also list the total H-He mass fractions and compare them to previous work in Table 2. We find that the total H-He mass fractions (0.11–0.23 for Uranus and 0.10–0.19 for Neptune) are consistent with previous studies with the exception of Militzer (2024), who obtained lower values. The derived rock-to-water ratios of our models (and selected previous studies) are listed in Table 2. We find that for both Uranus and Neptune, the rock-to-water ratio is poorly constrained and ranges between 0.04 and 3.92 for Uranus, and between 0.20 and 1.78 for Neptune. Overall, our models include solutions with rock-to-water ratios that are significantly higher compared to previous studies (up to 3.92 for Uranus and 1.78 for Neptune). At the same time, we show that solutions with low rock-to-water ratios (as low as 0.04 for Uranus and 0.20 for Neptune) are also valid. We can therefore conclude that given the current data, there is no clear reason to prefer water-rich or rock-rich interiors for Uranus and Neptune. Therefore, referring to Uranus and Neptune as ”ice giants” may be inappropriate since the bulk composition of the planets remain unclear (for example Helled & Fortney, 2020; Hofstadter et al., 2024). Constraining the rock-to-water ratio would require additional (and more accurate) measurements (for example Nimmo et al., 2024) or additional constraints from planetary formation models (for example Helled, 2025).
@@ 5.2 Magnetic field generation
Thanks to Voyager 2, we know that both Uranus and Neptune possess considerable magnetic fields (Connerney et al., 1987; Ness et al., 1989). The fields are multipolar and tilted significantly relative to the rotation axes of the planets (for example Jones, 2011). These observations should be considered when modelling the planetary interiors, as solutions should be consistent with dynamo generation. A common explanation for the magnetic fields of the planets is the presence of ionic water in a convective region (Redmer et al., 2011). For example, Stanley & Bloxham (2004, 2006) found that a thin convective and conductive shell in the outer planetary region can suppress large-scale dipolar modes and hence explain the observed multipolar fields. This occurs for both liquid (+stably stratified) and solid (+insulating) cores. And Gastine et al. (2012) showed that multipolar fields can be reproduced for an even wider range of scenarios.
Figure 7: Convective and stable regions in our models for Uranus and Neptune. Also shown are regions where ionic water is expected (see text for discussion).
@@ 5.3 Immiscibility of hydrogen, helium and water mixtures
Our models are based on the ideal mixing approximation (see Appendix B and Equation 10 for more details). This assumption can lead to differences in density of the order of a few percent (for example Darafeyeu et al., 2024). In particular, under this assumption immiscibility between different elements is not considered. Recently, Howard et al. (2025) presented miscibility results based on ab initio calculations of the hydrogen-water phase diagram. However, since our models include a mixture of hydrogen, helium, and water, a correction is needed. Using their Equation A.1, we can calculate the pressure-temperature line, below which hydrogen, helium, and water become immiscible. We employ:
@@
In the outermost convective layers, we tend to find higher H-He mass fractions for Uranus (0.62–0.73) compared to Neptune (0.25–0.49). This is consistent with the findings of Nettelmann et al. (2013); Cano Amoros et al. (2024); Tejada Arevalo (2025) and could indicate that Uranus indeed possesses more H-He in the outermost regions compared to Neptune. Furthermore, the outermost edge $r_{2}$ (see Table 3) of the magnetic dynamo regions tend to be slightly higher up for Neptune (0.78–0.92) compared to Uranus (0.69–0.74). This is consistent with the findings of Militzer (2024) and therefore also likely indicative of a dichotomy between Uranus and Neptune. Atmospheric probes sent to the atmospheres of both planets could test these predictions. Regarding the general bulk compositions (Figure 6) or the inferred rock-to-water ratios (Table 2), we remain agnostic about inferring concrete differences between Uranus and Neptune. We can only confirm that a wide range of options are possible for both planets.
We further note that the two planets have different measured heat fluxes (Pearl & Conrath, 1991) which is often explained with different structures such as gradients or boundary layers (for example Vazan & Helled, 2020; Scheibe et al., 2019, 2021) or giant impacts (for example Reinhardt et al., 2020). Our models do not include energy balance or flux calculations. If differences in structures are indeed the reason for the different heat fluxes, that would favour models U1 and U3 for Uranus and models N2 and N3 for Neptune. U1 and U3 contain extensive composition gradients where heat transport can not be facilitated efficiently via convection. As a consequence, comparatively little heat would be able to escape Uranus, while the large convection zones in models N2 and N3 would enable a more efficient heat transport and therefore a larger heat flux in comparison to Uranus. Interestingly, models U1 and U3 are rock-poor, while models N2 and N3 are rock-rich. While this could be indicative of a bulk composition dichotomy between Uranus and Neptune, we refrain from making any definite statements. For a robust conclusion, both more models and heat flux calculations (and measurements) would be required. We hence remain agnostic about this issue and hope to investigate this topic in detail in future work.
@@ 6 Discussion
This work introduces a novel method for modelling the internal structures of planetary interiors. By uniting the flexibility of empirical models with the physical validity of traditional approaches, we produce agnostic and self-consistent interior solutions. Our iterative algorithm allows for the convergence of randomly initialized density profiles. They converge toward solutions that simultaneously satisfy hydrostatic equilibrium, fit the gravity data, and respect constraints imposed by the employed EoS. We note that interior models can be further constrained by seismology Mankovich et al. (2025). We provide realistic Brunt-Väisälä frequencies necessary for interpreting oscillation modes in planetary interiors in Figures 8 and 10. Our models are well suited for comparison with asteroseismic observations by future missions. While this study represents a major step forward in planetary modelling, some limitations remain and future research is clearly desirable.
First, the EoS of materials at planetary conditions themselves may contain errors that can affect the results (More et al., 1988; Chabrier & Debras, 2021; Haldemann et al., 2020; Cozza et al., 2025). For example, the calculations in Haldemann et al. (2020) contain an error that affects the entropy, but not its derivatives (Aguichine et al., 2025). Tejada Arevalo (2025) found that correcting for this error produces qualitatively similar results. All EoSs of pure materials possess uncertainties and it is hence clear that the choice of the EoS affects the results. In Appendix D, we compare two different EoS for SiO2 to quantify how different their results are as an additional example. We find that the differences in density can be of the order of a few percent. Furthermore, the assumption of ideal mixing (Equation 10) can yet again lead to differences in density of the order of a few percent (for example Darafeyeu et al., 2024) and therefore affect the inferred composition. So even if the EoSs of the pure materials would serve as perfect constraints, a complexity arises from the fact that the mixtures are treated in an over-simplified manner. Finally, we only considered immisciblity for the hydrogen-helium-water mixture, but did not include other immiscibitilies (like a mixture of water and rock). Overall, we can therefore expect that the uncertainties associated with all of the aforementioned issues would also affect the inferred bulk compositions.
@@
Third, our models rely on the pressure-density relation provided in Hueso et al. (2020) for pressures below 100 bars. We rely on an atmospheric model because our approach is not well suited to handle the complex chemical processes occurring in planetary atmospheres such as condensation. Atmospheric models are generally more detailed and therefore more accurate in representing the conditions in the planetary atmosphere. Therefore, we rather not assume an atmospheric composition that could bias the results of the interior model. However, we do ensure that our models have density-pressure relations that are compatible with these more accurate atmospheric models (see Morf et al., 2024, for details). Although the used atmosphere model by Hueso et al. (2020) should be more accurate than our approach below 100 bars, it also includes uncertainties. To account for this, we allowed the initial temperature $T^{0}_{\text{EoS}}$ to deviate from the atmosphere model. Nonetheless, a different atmosphere model could lead to different results.
Fourth, we did not consider dynamical wind corrections to the measured gravitational moments (Kaspi et al., 2013; Soyuer et al., 2023). This can affect the density profile solution space and therefore our findings. However, Neuenschwander et al. (2024) investigated this effect for Uranus and suggested that wind corrections should not have a large influence on the inferred temperature and composition profiles. Since Uranus’ gravity data are significantly more constrained than Neptune’s, we expect that the same holds true for Neptune.

### v2

@@ 3 Uranus
Figure 3: Density, pressure, and temperature profiles for the four Uranus models. Top: Density versus normalised radius. For comparison, also shown are empirical solutions from Neuenschwander & Helled (2022) (grey area). The panel in the top right shows Uranus’ measured (French et al., 2024, black) and calculated (coloured) gravitational moments. We also show Uranus’ gravitational moments according to Jacobson & Park (2025) (grey) who find different results (COR vs COO) corresponding to different ring centre definitions. The coloured uncertainties depict estimated errors from the Theory of Figures. Bottom: Temperature versus pressure. Also shown is the predicted region of ionic water inferred by Redmer et al. (2011). The dashed lines show the demixing line, below which hydrogen, helium, and water are expected to become immiscible (Howard et al., 2025).
We present four Uranus models: U1, U2, U3, and U4. The inferred internal structures and compositions of the Uranus models are summarised in Figure 2. Models U1 and U3 correspond to water-rich interior solutions. Their respective rock-to-water ratios are low, just 0.04 and 0.08, consistent with Uranus being an ’ice giant’. These models also possess similar total H-He abundances (14% and 11% of Uranus’ mass, respectively). However, their convection and temperature profiles vary significantly. Model U1 is mostly convective, has four different convection zones and reaches a central temperature below 13,000 K. On the other hand, model U3 has two large stable regions. The innermost stable region covers more than half (54%) of Uranus’s radius and the central temperature is above 20,000 K.
@@
The density profiles and temperature-pressure profiles are presented in Figure 3. We find that the density profiles for models U2 and U4 are very similar. They both have a density discontinuity ($\Delta\rho>$ 1 g cm-3) at $\sim$ 75% of Uranus’ radius and reach central densities of $\sim$4.5 g cm-3. This density discontinuity around $3\cdot 10^{5}$ bar leads to a temperature jump that is 500 K higher for U2 compared to U4. Models U1 and U3 reach higher central densities (6.6 and 7.8 g cm-3, respectively) than U2 and U4. There are also no clear jumps in density in models U1 and U3. We further observe that all models leave the solution space discovered by Neuenschwander & Helled (2022) around a normalised radius of 0.8. This finding demonstrates how assumptions made by the modeller (such as the use of three consecutive polytropes in Neuenschwander & Helled (2022)) can result in biased outcomes and exclude parts of the solution space. In contrast, our randomised approach represents a notable improvement, expanding the scope of the solution space.
For comparison, we show the measured gravitational moments from Jacobson & Park (2025) (in grey) alongside those from French et al. (2024) (in black) in Figure 3. Jacobson & Park (2025) only report formal (not realistic) uncertainties for their two results based on precession rates of either the ring centres of opacity (COO) or the geometric ring midlines (COR). They conclude that their knowledge of the gravitational moments is not better compared to that postulated by French et al. (2024).
@@ 4 Neptune
We present four Neptune models N1, N2, N3 and N4. Figure 4 shows sketches of the inferred structure and bulk composition of our Neptune models. Models N1 and N4 are consistent with the notion of Neptune being an ’ice giant’. They have water-dominated interiors with rock-to-water ratios of 0.20 and 0.26, respectively, and H-He mass fractions of 12% and 10%, respectively. However, there are also significant differences. Model N1 has a central density higher than 18 g cm-3, while the central density of model N4 is only 6.2 g cm-3. Although both models include stable regions, they are more prevalent for model N1. For example, model N4 has a fully convective deep interior while the innermost region of model N1 is stable against convection and has a composition gradient. Model N4 initially has a larger temperature gradient compared to model N1, leading to a first appearance of ionic water at relatively larger normalised radii (0.89 vs 0.83) with higher temperatures (3100 K vs 2500 K).
@@
Models N2 and N3 represent Neptune models that are rock-dominated, supporting the idea of Neptune being a ’rock giant’ (for example Helled & Fortney, 2020; Teanby et al., 2020). Their rock-to-water ratios are found to be 1.78 and 1.72, with more than 50% of the planetary mass consisting of rocks. Although both models have large convective zones, the interior temperature profiles are rather different. Model N2 has a central temperature of 28,000 K, while the central temperature of model N3 is 12,000 K. Both of these inferred temperatures are significantly higher than the temperature inferred by standard adiabatic models which are of the order of 6000 K (for example Nettelmann et al., 2013). In addition, in model N3 H-He is found also in the planetary centre with 10 % of the composition by mass. The existence of H-He in the deep interior is consistent with formation models of Neptune with composition gradients (for example Valletta & Helled, 2022). The central region in model N2, on the other hand, is enriched with iron (18 % by mass).
@@
The inferred density profiles and pressure-temperature profiles are shown in Figure 5. For Neptune, we find four very unique density profiles. Model N1 has the highest central density of 18.4 g cm-3. The density in model N2 is below 3 g cm-3 in the outer half of the planet until it reaches a large density discontinuity ($\Delta\rho>3$ g cm-3). Models N3 and N4 have similar central densities (5.8 and 6.2 g cm-3, respectively), but differ with respect to the location of their density discontinuity. The discontinuities are located at a normalised radius of $\sim$0.7 for model N3 and $\sim$0.6 for model N2. We further find that models N1, N2, and N4 have a temperature jump larger than 500 K between pressures of 200 and 2000 bars. Also in the case of Neptune, there are significant differences in the density profiles in comparison to the solution space found by Neuenschwander & Helled (2022). This is observed at multiple locations (normalised radii of around 0.05, 0.4, and 0.75) with the differences being more notable compared to Uranus. Our randomised approach indeed covers a wider space of solutions compared to parametrised approaches.
@@ 5.1 Bulk composition
The inferred bulk compositions of our Uranus and Neptune models are summarised in Figure 6. We also list the total H-He mass fractions and compare them to previous work in Table 2. We find that the total H-He mass fractions (0.11–0.23 for Uranus and 0.10–0.19 for Neptune) are consistent with previous studies with the exception of Militzer (2024), who obtains lower values. The derived rock-to-water ratios of our models (and selected previous studies) are listed in Table 2. We find that for both Uranus and Neptune, the rock-to-water ratio is poorly constrained and ranges between 0.04 and 3.92 for Uranus, and between 0.20 and 1.78 for Neptune. Overall, our models include solutions with rock-to-water ratios that are significantly higher compared to previous studies (up to 3.92 for Uranus and 1.78 for Neptune). At the same time, we show that solutions with low rock-to-water ratios (as low as 0.04 for Uranus and 0.20 for Neptune) are also valid. We can therefore conclude that given the current data, there is no clear reason to prefer water-rich or rock-rich interiors for Uranus and Neptune. Therefore, referring to Uranus and Neptune as ’ice giants’ may be inappropriate since the bulk composition of the planets remain unclear (for example Helled & Fortney, 2020; Hofstadter et al., 2024). Constraining the rock-to-water ratio would require additional (and more accurate) measurements (for example Nimmo et al., 2024) or additional constraints from planetary formation models (for example Helled, 2025).
@@ 5.2 Magnetic field generation
Thanks to Voyager 2, we know that both Uranus and Neptune possess considerable magnetic fields (Connerney et al., 1987; Ness et al., 1989). The fields are multipolar and tilted significantly relative to the rotation axes of the planets (for example Jones, 2011). These observations are relevant when modelling the planetary interiors, as solutions have to be consistent with dynamo generation. A common explanation for the magnetic fields of the planets is the presence of ionic water in a convective region (Redmer et al., 2011). For example, Stanley & Bloxham (2004, 2006) found that a thin convective and conductive shell in the outer planetary region can suppress large-scale dipolar modes and hence explain the observed multipolar fields. This occurs for both liquid (+stably stratified) and solid (+insulating) cores. And Gastine et al. (2012) showed that multipolar fields can be reproduced for an even wider range of scenarios.
Figure 7: Convective and stable regions in our models for Uranus and Neptune. Also shown are regions where ionic water is expected.
@@ 5.3 Immiscibility of hydrogen, helium and water mixtures
Our models are based on the ideal mixing approximation (Appendix B and Equation 10 contain more details). This assumption can lead to differences in density of the order of a few percent (for example Darafeyeu et al., 2024). In particular, under this assumption immiscibility between different elements is not considered. Howard et al. (2025) provide miscibility results based on ab initio calculations of the hydrogen-water phase diagram. However, since our models include a mixture of hydrogen, helium, and water, a correction is needed. Using their Equation A.1, we can calculate the pressure-temperature line, below which hydrogen, helium, and water become immiscible. We employ:
@@ 5.4 Uranus versus Neptune
In the outermost convective layers, we tend to find higher H-He mass fractions for Uranus (0.62–0.73) compared to Neptune (0.25–0.49). This is consistent with the findings of Nettelmann et al. (2013); Cano Amoros et al. (2024); Tejada Arevalo (2025) and could indicate that Uranus indeed possesses more H-He in the outermost regions compared to Neptune. Furthermore, the outermost edge $r_{2}$ (Table 3) of the magnetic dynamo regions tend to be slightly higher up for Neptune (0.78–0.92) compared to Uranus (0.69–0.74). This is consistent with the findings of Militzer (2024) and therefore also likely indicative of a dichotomy between Uranus and Neptune. Atmospheric probes sent to the atmospheres of both planets could test these predictions. Regarding the general bulk compositions (Figure 6) or the inferred rock-to-water ratios (Table 2), we remain agnostic about inferring concrete differences between Uranus and Neptune. We can only confirm that a wide range of options are possible for both planets.
We further note that the two planets have different measured heat fluxes (Pearl & Conrath, 1991) which is often explained with different structures such as gradients or boundary layers (for example Vazan & Helled, 2020; Scheibe et al., 2019, 2021) or giant impacts (for example Reinhardt et al., 2020). Our models do not include energy balance or flux calculations. If differences in structures are indeed the reason for the different heat fluxes, that would favour models U1 and U3 for Uranus and models N2 and N3 for Neptune. Models U1 and U3 contain extensive composition gradients where heat transport can not be facilitated efficiently via convection. As a consequence, comparatively little heat would be able to escape Uranus, while the large convection zones in models N2 and N3 would enable a more efficient heat transport and therefore a larger heat flux in comparison to Uranus. Interestingly, models U1 and U3 are rock-poor, while models N2 and N3 are rock-rich. While this could be indicative of a bulk composition dichotomy between Uranus and Neptune, we refrain from making any definite statements. For a robust conclusion, both more models and heat flux calculations (and measurements) would be required. We hence remain agnostic about this issue and hope to investigate this topic in detail in future work.
@@ 6 Discussion
This work introduces a novel method for modelling the internal structures of planetary interiors. By uniting the flexibility of empirical models with the physical validity of traditional approaches, we produce agnostic and self-consistent interior solutions. Our iterative algorithm allows for the convergence of randomly initialised density profiles. They converge towards solutions that simultaneously satisfy hydrostatic equilibrium, fit the gravity data, and respect constraints imposed by the employed EoS. We note that interior models can be further constrained by seismology Mankovich et al. (2025). We provide realistic Brunt-Väisälä frequencies necessary for interpreting oscillation modes in planetary interiors in Figures 8 and 10. Our models are well suited for comparison with asteroseismic observations by future missions. While this study represents a major step forward in planetary modelling, some limitations remain and future research is clearly desirable.
First, the EoS of materials at planetary conditions themselves may contain errors that can affect the results (More et al., 1988; Chabrier & Debras, 2021; Haldemann et al., 2020; Cozza et al., 2025). For example, the calculations in Haldemann et al. (2020) contain an error that affects the entropy, but not its derivatives (Aguichine et al., 2025). Tejada Arevalo (2025) find that correcting for this error produces qualitatively similar results. All EoSs of pure materials possess uncertainties and it is hence clear that the choice of the EoS affects the results. In Appendix D, we compare two different EoS for SiO2 to quantify how different their results are as an additional example. We find that the differences in density can be of the order of a few percent. Furthermore, the assumption of ideal mixing (Equation 10) can yet again lead to differences in density of the order of a few percent (for example Darafeyeu et al., 2024) and therefore affect the inferred composition. So even if the EoSs of the pure materials would serve as perfect constraints, a complexity arises from the fact that the mixtures are treated in an over-simplified manner. Finally, we only considered immiscibility for the hydrogen-helium-water mixture, but did not include other immiscibilities (like a mixture of water and rock). Overall, we can therefore expect that the uncertainties associated with all of the aforementioned issues would also affect the inferred bulk compositions.
@@
Third, our models rely on the pressure-density relation provided in Hueso et al. (2020) for pressures below 100 bars. We rely on an atmospheric model because our approach is not well suited to handle the complex chemical processes occurring in planetary atmospheres such as condensation. Atmospheric models are generally more detailed and therefore more accurate in representing the conditions in the planetary atmosphere. Therefore, we rather not assume an atmospheric composition that could bias the results of the interior model. However, we do ensure that our models have density-pressure relations that are compatible with these more accurate atmospheric models (Morf et al., 2024, for details). Although the used atmosphere model by Hueso et al. (2020) should be more accurate than our approach below 100 bars, it also includes uncertainties. To account for this, we allowed the initial temperature $T^{0}_{\text{EoS}}$ to deviate from the atmosphere model. Nonetheless, a different atmosphere model could lead to different results.
Fourth, we did not consider dynamical wind corrections to the measured gravitational moments (Kaspi et al., 2013; Soyuer et al., 2023). This can affect the density profile solution space and therefore our findings. However, Neuenschwander et al. (2024) investigate this effect for Uranus and suggest that wind corrections should not have a large influence on the inferred temperature and composition profiles. Since Uranus’ gravity data are significantly more constrained than Neptune’s, we expect that the same holds true for Neptune.